solve 3x-1 divided by 5 = x-2
Answer by ewatrrr(24785)

3x-1 divided by 5 = x-2
3x-1 = 5(x-2)
3x-1= 5x-10
9 = 2x
9/2 = x = 4 1/2 = 4.5